Reports 1-1 of 1 Clear search Modify search
masahide.tamaki - 20:21 Tuesday 07 December 2021 (19117) Print this report
Health check (checking the power spectra of ITMX and ETMX -- payload part)

Related  work: klog19107

I checked the spectra of ITMX and ETMX payload (as for tower part, please see Ushiba-san's work).

In MN DAMP Y spectrum, there were three strange peaks around 0.7-1.1 Hz (surrouded by blue circle in Fig1).
*Solid line and Dashed line show the spectrum after/before finalization work.

With Ushibasan's advice, I checked the spectrum of BF and found that there were peaks at the same frequency (Fig2). So this may affects to peaks shown in Fig1.
Seeing the sensor spectrum, there are similarly peaks for H1, H2, H3 (Fig3).  So we can say L and T are well balanced.

Besides this (ETMX IM and MN, ITMX IM), there was no problem.

Images attached to this report
Comments to this report:
masahide.tamaki - 9:03 Wednesday 08 December 2021 (19123) Print this report

the frequency range in thir line is 0.07-0.11 Hz. Sorry.

Search Help